OpenStudy (anonymous):

Find the distance between points P(2, 2) and Q(7, 4) to the nearest tenth. A. 10.8 B. 4.5 C. 7 D. 5.4

OpenStudy (anonymous):

Use the distance formula

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

How did you put that up so fast >_>

OpenStudy (anonymous):

I got 7.5 ? I used the fomula you gave me

hartnn (hartnn):

x1 - x2 = = 5 y1-y2 = 4-2 = 2 is that what u git ?

hartnn (hartnn):

*got

OpenStudy (anonymous):

ohhh I got 5.38

hartnn (hartnn):

thats correct :) 5.4 rounded

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

welcome ^_^